A proton traveling at 107m/s Collides with a stationary particle and bounces back add 2×2 is collides with a stationary particle and bounces back at 2*10^6m/s. If the Particles moves forward at 3*10^6. Find the mass of the station iPad scree find the mass of the stationary Particle


1
Expert's answer
2021-11-25T10:14:31-0500

by momentum conservation:

"m_1v_1-m_1v_2=m_2v"

where m1, m2 are masses of proton and stationary particle,

v1, v2 is speed of proton before and after collision

v is speed of stationary particle after collision


then:


"m_2=\frac{m_1v_1-m_1v_2}{v}=\frac{1.67\cdot10^{27}(10^7-2\cdot10^6)}{3\cdot10^6}=6.68\cdot10^{-27}" kg